### v4

@@ Abstract
Consider a persuasion game where both the sender and receiver are ambiguity averse with maxmin expected utility (MEU) preferences and the sender can choose an ambiguous information structure. This paper analyzes the game in an ex-ante formulation: the sender first commits to an information structure, and then the receiver best responds by choosing an ex-ante message-contingent action plan. Under this formulation, I show it is never strictly beneficial for the sender to use an ambiguous information structure as opposed to a standard unambiguous one. This result is robust to (i) both players having (possibly heterogeneous) ambiguous beliefs over the states, and/or (ii) the receiver having non-MEU, uncertainty-averse preferences. However, it is not robust to the sender having non-MEU preferences.
@@ 1 Introduction
Without any additional information, the receiver strictly prefers New to Status Quo as the states are equally likely. Under Bayesian persuasion, the best the sender can achieve is to induce the receiver to choose Status Quo three quarters of the time. [^5] This is obtained by committing to a probabilistic information structure, a statistical experiment with two messages such that the receiver’s posterior given these two messages are 00 and $2/3$ in terms of the probability of $\omega_{2}$. The probabilities of sending these two messages are $1/4$ and $3/4$ respectively.
@@
Theorem 2 shows the conclusion of Theorem 1 continues to hold when (i) both players have (possibly heterogeneous) ambiguous beliefs over the states, and/or (ii) the receiver is an Uncertainty Averse decision maker as defined in Cerreia-Vioglio et al. (2011). However, one direction in which Theorem 1 is not robust, is regarding the assumption the sender’s preferences are MEU. In a follow-up paper, Cheng et al. (2024) demonstrate that a non-MEU ambiguity averse sender may obtain strict benefits from ambiguous communication, even when all other assumptions of Theorem 1 are preserved. While a full characterization of when and how the sender can benefit from ambiguous communication is provided in Cheng et al. (2024), the present paper includes a simple example (Example 2) to illustrate the possibility of such benefits for a non-MEU sender.
@@
While both perspectives are equally important, the ex-ante formulation may be more appropriate in certain contexts. Section 4 provides a detailed discussion of those settings. As a concrete illustration, Shishkin and Ortoleva (2023) conduct a lab experiment examining subject’s response to a decision problem and an ambiguous experiment identical to those described in Example 1. They find that a majority of the subjects choose the ex-ante optimal action. This behavior suggests that, when such subjects serve as receivers, a sender may fail to realize the gains from ambiguous persuasion identified by BLL, and could even do worse than using a standard statistical experiment—emphasizing the behavioral plausibility and practical relevance of the ex-ante formulation.
@@ 1.1 Related Literature
This paper contributes to the literature on Bayesian persuasion initiated by Kamenica and Gentzkow (2011), and more specifically, the line of research combining persuasion and ambiguity aversion (Beauchêne, Li and Li, 2019; Hedlund, Kauffeldt and Lammert, 2020; Nikzad, 2021; Cheng, 2022; Kosterina, 2022). While this paper shows that an MEU sender cannot benefit from ambiguous persuasion in the ex-ante formulation, in a follow-up paper, Cheng et al. (2024) characterize when and how a non-MEU sender may benefit. The ex-ante formulation adopted by the present paper, where the receiver chooses a message-contingent plan has also been used in other contexts, for example, the study of belief polarization (Baliga et al., 2013), information order (Li and Zhou, 2016, 2020; Wang, 2024), and incomplete information games (Hanany et al., 2020).
@@
There is a connection of this paper with papers on Bayesian persuasion when the players’ priors are heterogeneous (Alonso and Câmara, 2016; Laclau and Renou, 2017; Galperti, 2019). Given an ambiguous experiment, because the sender and receiver’s utility functions are different, the minimizer of their ex-ante payoffs may turn out to be different thus as if they have different priors. However, notice that this heterogeneity arises endogenously, while all the listed papers consider exogenously heterogeneous priors. Therefore, the results and findings of this paper cannot be obtained as implications of those papers. In addition, the main result of the paper is also shown to be robust to the players having (possibly heterogeneous) ambiguous beliefs over the states.
@@ 2 The Ex-Ante Formulation of the Persuasion Game
$&\max\limits_{\pi,f}&&U_{s}(\pi,f),\\ &\text{subject to }&&f\in\arg\max_{f\in\mathcal{F}}U_{r}(\pi,f).\end{aligned}$
@@
$&\max\limits_{\Pi,f}&&U^{MEU}_{s}(\Pi,f),\\ &\text{subject to }&&f\in\arg\max\limits_{f\in\mathcal{F}}U_{r}^{MEU}(\Pi,f).\end{aligned}$
@@ 3.2 On the Robustness of Theorem 1
    the players have (possibly heterogeneous) ambiguous beliefs over the states,
@@
since $\pi^{*}$ minimizes $U_{r}(\pi,bc)$ over all experiments in $\Pi$. Therefore, $bc$ can be induced when the sender commits to $\Pi$ [^10] For this $\Pi$, $ac$ is also a best response for the receiver. To address potential tie-breaking concerns, the sender can slightly perturb $\pi^{*}$ to generate posteriors $1/4-\eta$ and $3/4$ for an arbitrarily small $\eta>0$. This adjustment ensures that $bc$ becomes the receiver’s unique best response, while the sender’s payoff remains arbitrarily close to that under $\Pi$.
@@ Appendix A The Formalization and Proof of Theorem 2
This section formalizes the statements in Theorem 2 by introducing two alternative setups of the persuasion game, each corresponds to the relaxation of one of the original assumptions. The first setup allows the players to hold heterogeneous ambiguous beliefs over the states. The second setup permits the receiver to have more general Uncertainty Averse preferences. The case combining both relaxations is straightforward and thus is omitted.
A.1 Heterogeneous and Ambiguous Beliefs over States
Let $P_{s}$ and $P_{r}$ be convex and closed subsets of $\Delta(\Omega)$, representing the sender’s and receiver’s beliefs over the states, respectively. For simplicity, assume all measures in $P_{s}$ and $P_{r}$ have the same support. The sender’s program when choose among only statistical experiments is given by
$&\max\limits_{\pi,f}&&\min\limits_{p\in P_{s}}\sum_{\omega,m,a}f(m)(a)u_{s}(a,\omega)p(\omega)\pi(m|\omega),\\ &\text{subject to }&&f\in\arg\max_{f\in\mathcal{F}}\min\limits_{p\in P_{r}}\sum_{\omega,m,a}f(m)(a)u_{r}(a,\omega)p(\omega)\pi(m|\omega).\end{aligned}$ (A.1)
Similarly, when the sender is allowed to choose ambiguous experiments, the sender’s program is given by
$&\max\limits_{\Pi,f}&&\min\limits_{p\in P_{s}}\min\limits_{\pi\in\Pi}\sum_{\omega,m,a}f(m)(a)u_{s}(a,\omega)p(\omega)\pi(m|\omega),\\ &\text{subject to }&&f\in\arg\max\limits_{f\in\mathcal{F}}\min\limits_{p\in P_{r}}\min\limits_{\pi\in\Pi}\sum_{\omega,m,a}f(m)(a)u_{r}(a,\omega)p(\omega)\pi(m|\omega).\end{aligned}$ (A.2)
@@ A.1 Heterogeneous and Ambiguous Beliefs over States
$f\in\arg\max\limits_{f\in\mathcal{F}}\min\limits_{p\in P_{r}}\min\limits_{\pi\in\Pi}\sum_{\omega,m,a}f(m)(a)u_{r}(a,\omega)p(\omega)\pi(m|\omega).$
By the same argument, the minimax theorem applies and saddle points exist. Therefore, an optimal $f^{*}$ must be a saddle point such that it is optimal under some $p^{*}\in P_{r}$ and $\pi^{*}\in\Pi$. Then by substituting the saddle point into the receiver’s program, it follows that
$\max\limits_{f\in\mathcal{F}}\min\limits_{p\in P_{r}}\min\limits_{\pi\in\Pi}\sum_{\omega,m,a}f(m)(a)u_{r}(a,\omega)p(\omega)\pi(m|\omega)$
$=$ $\max\limits_{f\in\mathcal{F}}\min\limits_{p\in P_{r}}\sum_{\omega,m,a}f(m)(a)u_{r}(a,\omega)p(\omega)\pi^{*}(m|\omega)$
$=$ $\min\limits_{p\in P_{r}}\sum_{\omega,m,a}f^{*}(m)(a)u_{r}(a,\omega)p(\omega)\pi^{*}(m|\omega).$
That is, $f^{*}$ is also optimal for the receiver when the sender commits to $\pi^{*}$. Next,the sender’s payoff from committing to $\Pi$ satisfies
$U_{S}(\Pi,f^{*})$ $=\min\limits_{p\in P_{s}}\min\limits_{\pi\in\Pi}\sum_{\omega,m,a}f^{*}(m)(a)u_{s}(a,\omega)p(\omega)\pi(m|\omega)$
$\leq\min\limits_{p\in P_{s}}\sum_{\omega,m,a}f^{*}(m)(a)u_{s}(a,\omega)p(\omega)\pi^{*}(m|\omega)=U_{S}(\pi^{*},f^{*}).$
@@ A.2 Uncertainty Averse Receiver
Notice (i) - (iv) are standard conditions for a representation of Uncertainty Averse preferences. [^12] See Cerreia-Vioglio et al. (2011) for the required axioms and discussions. (v) is the additional requirement that the receiver views $C_{\Pi}$ as the set of relevant joint priors according to their Uncertainty Averse preferences. In other words, $G_{\Pi}$ is possibly finite only under those joint priors in $C_{\Pi}$. Notice that if the receiver has MEU preferences (as a special case of Uncertainty Averse preferences), this requirement simply reduces to the previous requirement on the receiver’s set of joint priors.
@@
$&\max\limits_{\Pi,f}&&U^{MEU}_{s}(\Pi,f),\\ &\text{subject to }&&f\in\arg\max\limits_{f\in\mathcal{F}}U^{UAP}_{r}(\Pi,f).\end{aligned}$

### v5

@@ Abstract
Consider a persuasion game where both the sender and receiver are ambiguity averse with maxmin expected utility (MEU) preferences and the sender can choose an ambiguous information structure. This paper analyzes the game in an ex-ante formulation: the sender first commits to an information structure, and then the receiver best responds by choosing an ex-ante message-contingent action plan. Under this formulation, I show it is never strictly beneficial for the sender to use an ambiguous information structure as opposed to a standard unambiguous one. This result is robust to (i) the players having heterogeneous beliefs over the states, and/or (ii) the receiver having non-MEU, uncertainty-averse preferences. However, it is not robust to the sender having non-MEU preferences.
@@ 1 Introduction
Without any additional information, the receiver strictly prefers New to Status Quo as the states are equally likely. Under Bayesian persuasion, the best the sender can achieve is to induce the receiver to choose Status Quo three quarters of the time. [^5] This is obtained by committing to a probabilistic information structure, a statistical experiment with two messages such that the receiver’s posterior given these two messages are 0 and $2/3$ in terms of the probability of $\omega_{2}$. The probabilities of sending these two messages are $1/4$ and $3/4$ respectively.
@@
Theorem 2 shows the conclusion of Theorem 1 continues to hold when (i) the players having heterogeneous beliefs over the states, and/or (ii) the receiver is an Uncertainty Averse decision maker as defined in Cerreia-Vioglio et al. (2011). However, one direction in which Theorem 1 is not robust, is regarding the assumption the sender’s preferences are MEU. In a follow-up paper, Cheng et al. (2024) demonstrate that a non-MEU ambiguity averse sender may obtain strict benefits from ambiguous communication, even when all other assumptions of Theorem 1 are preserved. While a full characterization of when and how the sender can benefit from ambiguous communication is provided in Cheng et al. (2024), the present paper includes a simple example (Example 2) to illustrate the possibility of such benefits for a non-MEU sender.
@@
While both perspectives are equally important, the ex-ante formulation may be more appropriate in certain contexts. A more detailed discussion appears in Section 4, but here I briefly outline two scenarios. First, when the sender communicates through a contingent contract and the receiver is legally required to commit at the ex-ante stage, the ex-ante formulation might be more suitable. Second, the receiver may adopt a dynamically consistent updating rule, as in Hanany and Klibanoff (2007, 2009), which ensures that their interim choices align with their ex-ante plan.
As a concrete illustration, Shishkin and Ortoleva (2023) conduct a lab experiment examining subject’s response to a decision problem and an ambiguous experiment identical to those described in Example 1. They find that a majority of the subjects choose the ex-ante optimal action. This behavior suggests that, when such subjects serve as receivers, a sender may fail to realize the gains from ambiguous persuasion identified by BLL, and could even do worse than using a standard statistical experiment—emphasizing the behavioral plausibility and practical relevance of the ex-ante formulation.
@@ 1.1 Related Literature
This paper contributes to the literature on Bayesian persuasion initiated by Kamenica and Gentzkow (2011), and more specifically, the line of research combining persuasion and ambiguity aversion (Beauchêne, Li and Li, 2019; Hedlund, Kauffeldt and Lammert, 2020; Nikzad, 2021; Cheng, 2022; Kosterina, 2022). While this paper shows that an MEU sender cannot benefit from ambiguous persuasion in the ex-ante formulation, in a follow-up paper, Cheng et al. (2024) characterize when and how a non-MEU sender may benefit. [^8] Cheng et al. (2024) extend the analysis to the setting in which both players have smooth ambiguity preferences (Klibanoff et al., 2005), of which MEU is a special case. The ex-ante formulation adopted by the present paper, where the receiver chooses a message-contingent plan has also been used in other contexts, for example, the study of belief polarization (Baliga et al., 2013), information order (Li and Zhou, 2016, 2020; Wang, 2024), and incomplete information games (Hanany et al., 2020).
@@
There is a connection of this paper with papers on Bayesian persuasion when the players’ priors are heterogeneous (Alonso and Câmara, 2016; Laclau and Renou, 2017; Galperti, 2019). Given an ambiguous experiment, because the sender and receiver’s utility functions are different, the minimizer of their ex-ante payoffs may turn out to be different thus as if they have different priors. However, notice that this heterogeneity arises endogenously, while all the listed papers consider exogenously heterogeneous priors. Therefore, the results and findings of this paper cannot be obtained as implications of those papers. In addition, the main result of the paper is also shown to be robust to the players having heterogeneous beliefs over the states.
@@ 2 The Ex-Ante Formulation of the Persuasion Game
$\begin{aligned} &\max\limits_{\pi,f}&&U_{s}(\pi,f),\\ &\text{subject to }&&f\in\arg\max_{f\in\mathcal{F}}U_{r}(\pi,f).\end{aligned}$
@@
$\begin{aligned} &\max\limits_{\Pi,f}&&U^{MEU}_{s}(\Pi,f),\\ &\text{subject to }&&f\in\arg\max\limits_{f\in\mathcal{F}}U_{r}^{MEU}(\Pi,f).\end{aligned}$
@@ 3.2 On the Robustness of Theorem 1
    the players have heterogeneous beliefs over the states,
@@
since $\pi^{*}$ minimizes $U_{r}(\pi,bc)$ over all experiments in $\Pi$. Therefore, $bc$ can be induced when the sender commits to $\Pi$. [^11] For this $\Pi$, $ac$ is also a best response for the receiver. To address potential tie-breaking concerns, the sender can slightly perturb $\pi^{*}$ to generate posteriors $1/4-\eta$ and $3/4$ for an arbitrarily small $\eta>0$. This adjustment ensures that $bc$ becomes the receiver’s unique best response, while the sender’s payoff remains arbitrarily close to that under $\Pi$.
@@ Appendix A The Formalization and Proof of Theorem 2
This section formalizes the statements in Theorem 2 by introducing two alternative setups of the persuasion game, each corresponds to the relaxation of one of the original assumptions. The first setup allows the players to hold heterogeneous beliefs over the states. The second setup permits the receiver to have more general Uncertainty Averse preferences. The case combining both relaxations is straightforward and thus is omitted.
A.1 Heterogeneous Beliefs over States
Let $p_{s}$ and $p_{r}$ denote the sender and receiver’s priors over the states, respectively. The sender’s program when choosing among only statistical experiments is given by
$\begin{aligned} &\max\limits_{\pi,f}&&\sum_{\omega,m,a}f(m)(a)u_{s}(a,\omega)p_{s}(\omega)\pi(m|\omega),\\ &\text{subject to }&&f\in\arg\max_{f\in\mathcal{F}}\sum_{\omega,m,a}f(m)(a)u_{r}(a,\omega)p_{r}(\omega)\pi(m|\omega).\end{aligned}$ (A.1)
Similarly, when the sender is allowed to choose ambiguous experiments, the sender’s program becomes
$\begin{aligned} &\max\limits_{\Pi,f}&&\min\limits_{\pi\in\Pi}\sum_{\omega,m,a}f(m)(a)u_{s}(a,\omega)p_{s}(\omega)\pi(m|\omega),\\ &\text{subject to }&&f\in\arg\max\limits_{f\in\mathcal{F}}\min\limits_{\pi\in\Pi}\sum_{\omega,m,a}f(m)(a)u_{r}(a,\omega)p_{r}(\omega)\pi(m|\omega).\end{aligned}$ (A.2)
@@ A.1 Heterogeneous Beliefs over States
$f\in\arg\max\limits_{f\in\mathcal{F}}\min\limits_{\pi\in\Pi}\sum_{\omega,m,a}f(m)(a)u_{r}(a,\omega)p_{r}(\omega)\pi(m|\omega).$
By the same argument, the minimax theorem applies and saddle points exist. Therefore, an optimal $f^{*}$ must be a saddle point such that it is optimal under some $\pi^{*}\in\Pi$. Next,the sender’s payoff from committing to $\Pi$ satisfies
$U_{S}(\Pi,f^{*})$ $=\min\limits_{\pi\in\Pi}\sum_{\omega,m,a}f^{*}(m)(a)u_{s}(a,\omega)p_{s}(\omega)\pi(m|\omega)$
$\leq\sum_{\omega,m,a}f^{*}(m)(a)u_{s}(a,\omega)p_{s}(\omega)\pi^{*}(m|\omega)=U_{S}(\pi^{*},f^{*}).$
@@ A.2 Uncertainty Averse Receiver
Notice (i) - (iv) are standard conditions for a representation of Uncertainty Averse preferences. [^13] See Cerreia-Vioglio et al. (2011) for the required axioms and discussions. (v) is the additional requirement that the receiver views $C_{\Pi}$ as the set of relevant joint priors according to their Uncertainty Averse preferences. In other words, $G_{\Pi}$ is possibly finite only under those joint priors in $C_{\Pi}$. Notice that if the receiver has MEU preferences (as a special case of Uncertainty Averse preferences), this requirement simply reduces to the previous requirement on the receiver’s set of joint priors. [^14] Also see Section 5.2.2 in Cerreia-Vioglio et al. (2011) for the discussion in the case of smooth ambiguity preferences (Klibanoff et al., 2005).
@@
$\begin{aligned} &\max\limits_{\Pi,f}&&U^{MEU}_{s}(\Pi,f),\\ &\text{subject to }&&f\in\arg\max\limits_{f\in\mathcal{F}}U^{UAP}_{r}(\Pi,f).\end{aligned}$